OpenStudy (anonymous):

How do you find the domain for (foh)(x)= sqrt4-(x)^2 ?

OpenStudy (anonymous):

does the square root goo all the way through the equation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

alright so this makes it much easier. Since your square root goes through the whole equation, it means the value inside the square root cannot be negative. Given the equation \[f(x) = \sqrt{4 - x^2}\]

OpenStudy (anonymous):

x has to be a value between -2 and 2, since it cannot result in a value greater than 4 after doing the exponential operation. Therefore -2<x<2 is your domain.

OpenStudy (anonymous):

Thanks ! Would i do the same thing for \[\sqrt{(4-x)^2}\]

OpenStudy (anonymous):

No that is a set of all real numbers. Because even if x is greater than 4, squaring the negative would result in a positive value which can be square rooted. Say for instance x = 5 \[f(x) = \sqrt{(4 - x)^2}\] \[f(5) = \sqrt{(4-5)^2}\] \[f(5) = \sqrt{(4 -5)^2}\] \[f(5) = \sqrt{(-1)^2}\] \[f(5) = \sqrt{(1)}\] \[f(5) = 1

OpenStudy (anonymous):

\[f(5)\] ********

OpenStudy (anonymous):

Oh okay thank you :D!

OpenStudy (anonymous):

no problem and nice pic :) :P

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

I have one last question.. lol

OpenStudy (anonymous):

lol sure ask away :)

OpenStudy (anonymous):

Okay thanks lol Given: f(x)=\[\sqrt{4-x} \] and h(x)= x^2 Find g, the inverse of f. GIve the domain and range for both f and g, sketch the graphs of f and g on the same set of axes. Even if you could just help me with the domain and range that'd be awesome lol

OpenStudy (anonymous):

IM SOO SORRY, i got carried away by my own hmwk

OpenStudy (anonymous):

i havent got to my inverse hmwk yet, so i may not be the best source for the answer. :( Sorry once again :/